OpenStudy (anonymous):

Write log12144 = 2 in exponential form.

OpenStudy (campbell_st):

this is \[\log_{12} 144 = \log_{12} 12^2 = 2\log_{12} 12 = 2\times1 = 2 \] of simply \[12^{\log_{12} 144} = 12^2 .... 12^{\log_{12} 144} = 144\]

OpenStudy (anonymous):

thank you